OpenStudy (anonymous):

plz help gold medal will be rewardeBeth used zero point five pounds of sugar to empty a 5 pound bag. Beth wants to know how many pounds of sugar were previously used. Which of the following is a verbal model of this problem? sugar used by Beth plus total sugar originally in the bag equals sugar previously used sugar used by Beth minus sugar previously used equals total sugar originally in the bag total sugar originally in the bag plus sugar previously used equals sugar used by Beth sugar previously used plus sugar used by Beth equals total sugar originally in the bag

OpenStudy (anonymous):

Think of it logically, if someone else used your sugar, and you used the rest of it so there is no more left, your answer would be?

OpenStudy (anonymous):

Well you know she used 0.5 lbs of sugar and some amount previously so you can say, x + 0.5lbs = 5lbs. So This means answer D, because its sugar previously used plus sugar used by Beth equals the total sugar originally in the bag.

OpenStudy (anonymous):

0

OpenStudy (anonymous):

is that right

OpenStudy (anonymous):

sugar used by Beth plus total sugar originally in the bag equals sugar previously used

OpenStudy (anonymous):

Well that would be 0.5+5=5.5 which is more than originally in the bag, so no.

OpenStudy (anonymous):

total sugar originally in the bag plus sugar previously used equals sugar used by Beth

OpenStudy (anonymous):

ugar previously used plus sugar used by Beth equals total sugar originally in the bag

OpenStudy (anonymous):

sugar used by Beth minus sugar previously used equals total sugar originally in the bag

OpenStudy (anonymous):

1,2,3, or 4

OpenStudy (anonymous):

Ok lets try this find the previous amount used by finding x in (x+0.5=5). What is x?

OpenStudy (anonymous):

uhhhhhh

OpenStudy (anonymous):

im russian to the fridge

OpenStudy (anonymous):

Ok what is 5-0.5

OpenStudy (anonymous):

thats easy

OpenStudy (anonymous):

Ok what is it?

OpenStudy (anonymous):

4.5

OpenStudy (anonymous):

Ok so that value there is sugar previously used. So now if you have sugar previously used plus sugar used by Beth you get what?

OpenStudy (anonymous):

1964

OpenStudy (anonymous):

Let me know when you want to figure this out